# Strong asymptotic giant branch stars’ spectral features in distant quiescent galaxies: Impact on galaxy evolution

### v1

@@ Abstract
Age-dating and weighting stellar populations in galaxies at various cosmic epochs are essential steps to study galaxy formation through cosmic times. Evolutionary population synthesis models with different input physics are used towards this aim. In particular, the contribution from the thermally pulsing asymptotic-giant-branch (TP-AGB) stellar phase, which peaks for intermediate-age 0.6–2 Gyr systems, has been debated upon for decades. Here we report the detection of strong cool star signatures in the rest-frame near-infrared spectra of three young (1 Gyr), massive () quiescent galaxies at large look-back time, z=1–2, using JWST/NIRSpec. The co-existence of oxygen- and carbon-type absorption features, spectral edges and features from rare species such as Vanadium, and possibly Zirconium, reveal a strong contribution from TP-AGB stars. Population synthesis models with significant TP-AGB contribution reproduce the observations considerably better than those with weak TP-AGB, which are those commonly used. These findings call for revisions of published stellar population fitting results, pointing to lower masses and younger ages, with additional implications on cosmic dust production and chemical enrichment. These results will stimulate new generations of improved models informed by these and future observations.
@@ 1 Introduction
The TP-AGB – an advanced evolutionary stage of intermediate-mass stars (1-10) – is a potentially essential component of stellar population models. The study of the TP-AGB contribution to the light of stellar population has a long history. The first estimate goes back over 30 years [1] and was based on the fuel consumption theorem, thus predicting that the TP-AGB contribution to the bolometric light of a simple stellar population (SSP) could approach or even exceed at intermediate ages from Myr to a few Gyr. Subsequent models estimated a contribution to up to 80% of the integrated near-infrared (NIR) emission of stellar populations with ages in the range of 0.2 2 Gyr [2, 3, 4, 5, 6].
@@
However, the actual energetic contribution and spectral composition of this phase are highly controversial, with different models adopting distinctively different prescriptions for both the energetics and the stellar spectra during this phase [7, 3, 8, 9]. These uncertainties stem from two complications typical of the TP-AGB phase, namely its double-shell burning regime leading to instabilities and ‘thermal pulses’ occurring on short timescales alongside a strong mass-loss, both requiring calibration of free parameters such as convection, mixing and mass loss with observational data [10, 11]. As a consequence, stellar evolutionary computations, which are the backbone of evolutionary population synthesis models, typically terminate at the end of the Early Asymptotic Giant Branch (E-AGB) phase, and the inclusion of the TP-AGB in integrated models is performed separately either using specific TP-AGB tracks [8, 4] or fuel-consumption-theorem estimated TP-AGB energetics [3].
@@
Historically, the early Maraston models (i.e. M05 [3]) based on the fuel consumption theorem [12] pioneered the incorporation of the TP-AGB phase into population synthesis models by calibrating the energetics of TP-AGB tracks with observed photometric data and adopting observed oxygen-rich (O-rich) and carbon-rich (C-rich) spectra [13, 14]. These M05 models, forecasting the onset of the TP-AGB phase as a sharp transition (e.g., a rapid increase in the TP-AGB bolometric and near-IR contribution to the energetics of stellar populations, see Figure 3 in Ref. [2]) around 0.2 Gyr and lasting up to 2 Gyr, are referred to as “TP-AGB-heavy”. A similar modelling approach is found in the Conroy models (i.e., C09 [9]), but with two modifications, namely a substantial reduction of the energetics and a shift towards older ages (1-3 Gyr) for the relevant population epoch, which dilute the TP-AGB ‘phase transition’. Therefore, the C09 models are referred to as “TP-AGB-light”. Following the availability of new data for the calibration and adopting some of the C09 arguments, the calibration of the TP-AGB contribution as a function of age was revised using Magellanic Clouds globular clusters data [10], based on which an update of the M05 models was presented (so-called M13). The M13 models have a reduced TP-AGB fuel consumption and a slightly older onset age (600 Myr) with respect to M05, but they still forecast a sharp transition for the TP-AGB onset and a sizable fuel consumption. The M13 models are referred to as “TP-AGB-mild”. On a different route, the widely used Bruzual & Charlot models (i.e., BC03 [7]), based on isochrone synthesis of stellar evolutionary tracks [15], adopt different energetics and spectra for the TP-AGB phase. Since this results in a negligible TP-AGB contribution to the integrated spectrum, the BC03 models are referred to as “TP-AGB-poor”.
@@
A number of studies have attempted to observationally constrain the contribution of the TP-AGB phase by fitting spectra and photometric spectral energy distributions (SEDs) of different kinds of galaxies (post-starbursts, Seyferts, spirals, and passive) over a range of redshifts, but with conflicting results [16, 17, 18, 19, 20, 21]. It was early suggested (M05) that high-redshift galaxies with low star formation (‘quiescent’) are the ideal laboratories for this test, as their mass-weighted ages (1 Gyr) correspond to the expected maximum possible contribution of the TP-AGB stars to the rest-frame NIR spectrum of such galaxies. High quality observations of NIR galaxy spectra dominated by TP-AGB stars are still missing: ground-based spectra below 2.3m are impacted by atmospheric emission and absorption, and HST NIR grism spectra lack sufficient sensitivity. The JWST has now the wavelength coverage and sensitivity to finally settle this issue.
@@ 2 Results
2.1 Selection of quiescent candidates
@@
We selected 1084 quiescent galaxy candidates at using a catalog [22] in the EGS CANDELS field [23, 24], based on the UVJ diagram [25] and SED modeling (adopting exponentially declining star formation history, namely SFH and requiring ). The limit imposed on the sample ensures that the age-sensitive spectral region around rest-frame 4000Å break is covered by JWST/NIRSpec with good sensitivity. We cross-matched these phometrically-selected candidates with the 1491 galaxies with secure spectroscopic redshift observed with JWST/NIRSpec during the Cosmic Evolution Early Release Science (CEERS) program [26, 27] and during a subsequent NIRSpec follow-up program (DD-2750) [28]. This resulted in a spectroscopic sample of 31 galaxies, for which we visually inspected the spectra. Some spectra have low signal-to-noise ratios (SNR) because the galaxies are too faint or were poorly centered in the NIRSpec’s micro-shutter apertures. Other spectra exhibit strong emission lines consistent with either active star formation or an active galactic nucleus (AGN), or that have redshifts that are too high () to provide adequate coverage of the relevant rest-frame NIR wavelengths. We retained three galaxies, IDs = 8595, 9025, and D36123 with redshifts at , whose spectra clearly show the D4000 break, no obvious emission lines, and also have relatively high signal-to-noise ratio (SNR). This is a small yet representative sample fulfilling these criteria.
2.2 The spectrum of D36123
@@
Galaxy D36123, the primary target in this work, exhibits an exceptionally high-quality PRISM spectrum with an average of SNR 187 per spectral pixel (Fig.1; plotted as a double histogram to show the noise). Its relatively bright flux (mag=21.4 at 3.6m), steep radial surface brightness profile (Sérsic 4, see Table 3), longer exposure time and accurate centering into the micro-shutters, which minimizes the effect of path loss and aperture correction (see Methods), are the reasons for the remarkable quality of this spectrum.
@@
A large number of features are clearly detected at rest-frame wavelengths from 0.3 to 2.0 m, for which we propose identifications (see labels in Fig.1). Besides common absorption features at Å typical of quiescent galaxies, such as Ca H+K, H, H+G4300 and absorption from heavy elements (such as Mg, Ca, and Fe), which have been observed in quiescent galaxies’ spectra at high redshift [29], the rest-frame NIR spectrum exhibits numerous absorption features and band edges. The relatively low spectral resolution (R30–300) of the NIRSpec PRISM helps in the detection of these broad features. The proposed identifications are obtained by comparing our spectrum with empirical libraries of C-rich [14] and O-rich [13] TP-AGB stars (see Fig.2). These observed spectral features are related to diverse stellar types in the AGB phase, i.e. E-AGB and TP-AGB [30, 31, 32, 33]. The deep CN edges in Fig.2a (most striking is the CN1.1m edge, see Fig.1, where the continuum suddenly drops down by 15% in the observed spectrum) are particularly enhanced in rich N-type carbon stars [30, 31, 32], that are produced by the third dredge-up and therefore demonstrate the presence of a contribution by TP-AGB stars. On the other hand, the TiO and VO absorption bands (see Fig.2b) are produced in the spectra of M-type stars in either the RGB or AGB phases [31]. However, the best-fit spectrum of BC03 models, that includes only the RGB contribution, is almost featureless in the NIR (Fig. 3) and we conclude that also the oxygen-rich features in the spectrum of D36123 are due to TP-AGB stars. This is the first time that multiple TiO absorptions concomitant with and features are identified in an observed galaxy spectrum, pointing to the simultaneous presence of O-rich and C-rich cool stars. This is naturally expected in presence of a broad distribution of stellar metallicities, which must exist in massive galaxies, such that the metal-poor stars evolve preferentially into carbon-rich TP-AGB stars, and the metal-rich ones into oxygen-rich (M Type) TP-AGB stars. Since V and Zr are only produced by s-process nucleosynthesis, those VO and ZrO features (see Fig.2b) are expected to be detectable only in the coolest TP-AGB stars [31, 33], demonstrating the prominence of the TP-AGB phase in the light emitted by this quiescent galaxy. Note also the broad depressions due to water, again indicating the presence of O-rich M-type stars.
@@
We then performed spectral fitting on the observed spectrum of D36123, employing four models – BC03, C09, M13, and M05 (spanning the full range from TP-AGB poor to heavy, see Fig.3 and Table 1). We used delayed exponentially declining SFH with a large variety of timescales and stellar ages, interpolated over a grid of metallicities, and allowed for dust extinction. In convolving the models to match the resolution of observed spectra, we considered the instrumental resolution, the wavelength sampling of the spectral template models, and the estimated stellar velocity dispersion of the galaxy (see Methods). We find that our model fitting results are robust against (1) more complex star-formation histories allowing for bursts and sudden truncations; (2) linear combinations of models with different ages and metallicities (which is equivalent to non-parametric SFH); (3) alternative attenuation curves or without considering the reddening by dust; (4) variations to the aperture correction recipe; (5) modifications of the spectral extraction apertures over 3 or 5 pixels; (6) modelling of emission lines. The minimum reduced from each model is barely affected when these variations are considered.
@@
The “TP-AGB-mild/heavy” (i.e., M13 and M05) models can produce the largest fraction of the detected features, while the “TP-AGB-poor/light” predict much weaker spectral features in the NIR, as Fig.3 illustrates. Among the four models, the TP-AGB-mild model, M13, provides by far the best fit. Additionally, M13 also identifies the majority of TP-AGB features (i.e., four TiO absorptions and most CN edges). However, there are strong spectral features that are clearly not present (or much weaker than in the observed spectrum) even in the M13 best-fit spectrum, like several TiO and C absorptions. The strong CN1.1m spectral edge is only present in the M13/M05 models. Interestingly, the MgII-band at m, part of Lick set of indexes [34] is much weaker in all models than in the data, pointing to a strong contribution from TiO bands all the way to 0.5m, as expected for M-type stars (and, possibly, to [/Fe]-enhancement [35]). The M05 shows the worst overall fit due to clear discrepancy at (see Fig.10 in Methods), but still displays significantly more numerous matching TP-AGB features in the NIR than the C09 and BC03 models (see Fig.3), noticeably the long-sought C absorption around 1.75m. Moreover, when fitting the spectrum over the restricted rest-frame wavelength range Å all models converge to a mass-weighted age of about 1 Gyr, which is indeed the age where TP-AGB features are expected to be at peak [3, 10]. Therefore, we can conclude that it is the treatment of the TP-AGB in the models that most affects the goodness of fit in our quiescent galaxies. Even for the best-fit model M13, the minimum reduced is large, indicating that also this model is lacking some stellar ingredients that are present in the galaxy, and are needed.
@@
Some of the observed features are also found in the coolest phase of (M-type) red-giant-branch (RGB) stars and in the red-supergiant (RSG) phase of massive stars, with ages around 10 Myr [13, 30, 33]. We could safely exclude that RSG stars contribute substantially due to the inferred mass-weighted age (1 Gyr) and negligible residual star formation rate, when fitting only the optical spectrum (see Table 1). This is suggestive of a negligible contribution from short-lived massive stars. Instead, M-type RGB stars certainly contribute to the NIR spectrum at these ages, to a degree that is, however, model-dependant because the onset and development of the RGB phase depends on the overshooting assumed in stellar tracks [36, 3]. The full RGB contribution to the total light in the K-band around 1 Gyr is according to M05 models and for BC03-type models and the upper part of the RGB where M-giants form is a further fraction of this. Therefore, while contributing, M-type RGB stars do not dominate the spectrum. We verified this point by comparing our observed spectrum to a further stellar population model including empirical spectra for the RGB (and other phases; E-MILES [37]), but not including a treatment of the TP-AGB phase. This model cannot reproduce more features than the BC03/C09 models.
@@
The best-fitting physical parameters presented in Table 1 exemplify the impact of the chosen model on the derived properties. For instance, comparing results between BC03 and M13 in the full rest-frame spectral range, there is only 30% discrepancy in stellar mass, but the impact on ages is strong with 1 Gyr (the M13 age is almost three times lower than the BC03 age – all reported age measurements in this work are mass-weighted), despite the higher metallicity preferred by the BC03 fit. These age differences affect the derived formation epochs and evolutionary timescales of galaxies. The BC03 best-fit age of 1.7 Gyr is consistent with a RGB contribution of about for those models (see Figure 2 in Ref.[38]), which likely contributes to some of the observed features and NIR flux. However, this fitted age is still too young for the contribution from red giant stars to be dominant, further supporting the arguments described above.

### v2

@@ Abstract
Age-dating and weighting stellar populations in galaxies are essential steps to study galaxy formation through cosmic times. Evolutionary population synthesis models with different input physics are used towards this aim. In particular, the contribution from the thermally pulsing asymptotic giant branch (TP-AGB) stellar phase, which peaks for intermediate-age 0.6–2 Gyr systems, has been debated upon for decades. Here we report the detection of strong cool-star signatures in the rest-frame near-infrared spectra of three young (1 Gyr), massive () quiescent galaxies at large look-back time, z=1–2, using JWST/NIRSpec. The co-existence of oxygen- and carbon-type absorption features, spectral edges and features from rare species such as Vanadium, and possibly Zirconium, reveal a strong contribution from TP-AGB stars. Population synthesis models with significant TP-AGB contribution reproduce the observations better than those with weak TP-AGB, which are commonly used. These findings call for revisions of published stellar population fitting results, pointing to lower masses and younger ages, with additional implications on cosmic dust production and chemical enrichment. New generations of improved models are needed, informed by these and future observations.
@@ 1 Introduction
The TP-AGB – an advanced evolutionary stage of low- and intermediate-mass stars (0.8-10) – is a potentially essential component of stellar population models. Quantifying the TP-AGB contribution to the light of stellar populations has a long history. The first estimate goes back over 30 years[1], which was based on the fuel consumption theorem[2], predicted that the TP-AGB contribution to the bolometric light of a simple stellar population (SSP) could approach or even exceed 50% at intermediate ages. Subsequent models estimated a contribution of up to 80% of the integrated near-infrared (NIR) emission of stellar populations with ages in the range of 0.2 2 Gyr[3, 4, 5, 6, 7].
@@
The actual energetic contribution and spectral composition of this phase, however, remain highly controversial, with different models adopting distinctively different prescriptions for both the energetics and the stellar spectra[8, 4, 9, 10]. These uncertainties stem from two complications typical of the TP-AGB phase, namely its double-shell burning regime leading to instabilities and ‘thermal pulses’ occurring on short timescales alongside a strong mass-loss, both requiring calibration of free parameters such as convection, mixing and mass loss with observations[11, 12]. As a consequence, stellar evolutionary computations, which are the backbone of evolutionary population synthesis models, typically terminate at the end of the Early Asymptotic Giant Branch (E-AGB) phase, and the inclusion of the TP-AGB in integrated models is performed separately either using specific TP-AGB tracks[9, 5] or fuel-consumption-theorem estimated TP-AGB energetics[4].
@@
Historically, the early Maraston models (i.e. M05[4]) based on the fuel consumption theorem[2] pioneered the incorporation of the TP-AGB phase into population synthesis models by calibrating the energetics of TP-AGB tracks with observed photometric data and adopting observed oxygen-rich (O-rich) and carbon-rich (C-rich) spectra[13, 14]. These M05 models, forecasting the onset of the TP-AGB phase as a sharp transition (e.g., a rapid increase in the TP-AGB bolometric and NIR contribution to the energetics of stellar populations[3]) around 0.2 Gyr and lasting up to 2 Gyr, are referred to as “TP-AGB-heavy”. A similar approach is found in the Conroy models (i.e., C09[10]), but with two modifications, namely a substantial reduction of the energetics and a shift towards older ages (1-3 Gyr) for the relevant population epoch, which dilute the TP-AGB ‘phase transition’. The C09 models are referred to as “TP-AGB-light”. Following the availability of new data for the calibration and adopting some of the C09 arguments, the calibration of the TP-AGB contribution as a function of age was revised using Magellanic Clouds globular clusters data[11], and an update of the M05 models was presented (M13 models; referred to as “TP-AGB-mild”) having a reduced TP-AGB fuel consumption and a slightly older onset age (600 Myr) with respect to M05, but still predicting a sharp transition for the TP-AGB onset and a sizable fuel consumption. Finally, the widely used Bruzual & Charlot models (i.e., BC03[8]), based on isochrone synthesis of stellar evolutionary tracks[15], adopt different energetics and spectra for the TP-AGB phase. Since this results in a negligible TP-AGB contribution to the integrated spectrum, the BC03 models are referred to as “TP-AGB-poor”.
@@
Many studies have attempted to observationally constrain the contribution of the TP-AGB phase by fitting spectra and photometric spectral energy distributions (SEDs) of different types of galaxies (post-starbursts, Seyferts, spirals, and passive) over a range of redshifts, with conflicting results[16, 17, 18, 19]. Quiescent high-redshift galaxies (mass-weighted ages 1 Gyr producing peak TP-AGB-stars contribution to the NIR) are ideal laboratories but until now only imaging data were available at high redshift. High-quality NIR galaxy spectra dominated by TP-AGB stars are still missing: ground-based spectra below 2.3m are impacted by atmospheric emission and absorption, and HST NIR grism spectra lack sufficient sensitivity. JWST has now the wavelength coverage and sensitivity to finally settle this issue.
@@ 2 Results
2.1 Selection of quiescent galaxies
@@
We searched for distant quiescent galaxies serendipitously included (as fillers) in the Cosmic Evolution Early Release Science (CEERS) JWST NIRSpec program[20, 21] and during a subsequent NIRSpec follow-up program (DD-2750)[22]. We started identifying 1084 quiescent galaxy candidates at 1 in the EGS CANDELS[23, 24] catalog[25], based on the UVJ diagram[26] and SED modelling (adopting exponentially declining star formation history, namely SFH and requiring 1.0). The 1 limit ensures that the age-sensitive spectral region around rest-frame 4000Å break is covered by JWST/NIRSpec with good sensitivity. We cross-matched these photometric candidates with the 1491 galaxies already with secure NIRSpec spectroscopic redshift. This resulted in 31 galaxies whose spectra we visually inspected. Some spectra have low signal-to-noise ratios (SNR) because the galaxies are faint or poorly centered in the NIRSpec’s micro-shutter apertures. Other spectra exhibit strong emission lines consistent with either active star formation or an active galactic nucleus (AGN). A few have too high redshifts (3) to provide adequate coverage of the relevant rest-frame NIR wavelengths. We retained three galaxies, IDs = 8595, 9025, and D36123 with redshifts at 12, whose spectra clearly show the D4000 break, no obvious emission lines, and have relatively high SNR.
2.2 The primary spectrum of Galaxy D36123
@@
Galaxy D36123 exhibits an exceptionally high-quality PRISM spectrum with an average of SNR187 per spectral pixel (Fig.1; plotted as a double histogram showing the noise). The relatively bright flux (magAB=21.4 at 3.6m), steep radial surface brightness profile (Sérsic 4, see Extended Data Fig.1), longer exposure time and accurate centering into the micro-shutters, which minimizes the effect of path loss and aperture correction (see Methods), are the reasons for the remarkable quality of this spectrum.
@@
A large number of features are clearly detected at rest-frame wavelengths from 0.3 to 2.0 m, for which we propose identifications (see labels in Fig.1 and Supplementary Table 1). Besides common absorption features at 5500Å typical of quiescent galaxies, such as Ca H+K, H, H+G4300 and absorption from heavy elements such as Mg, Ca, and Fe, which have been observed in quiescent galaxies at high redshift[27], the spectrum exhibits numerous absorption features and band edges. The relatively low spectral resolution (R30-300) of the NIRSpec PRISM helps in the detection of these broad features. The proposed identifications are obtained by comparison with empirical libraries of C-rich[14] and O-rich[13] TP-AGB stars (see Fig.2a,b). These features are related to diverse stellar types in the AGB stellar phases, i.e. E-AGB and TP-AGB[28, 29, 30, 31]. The deep CN edges in Fig.2a (most striking is the CN1.1m edge, see Fig.1, where the continuum suddenly drops down by 15%) are particularly enhanced in N-type carbon stars[28, 29, 30], which are produced by the third dredge-up and therefore demonstrate the presence of the contribution by TP-AGB stars (see CN1.1 index in Fig.3d,f). On the other hand, the TiO and VO absorption bands (Fig.2b) are produced in cool M-type stars in either the red-giant-branch (RGB) or AGB phases[29], but the equivalent width (EW) measured for specific features at 0.5-0.53 and 0.9-1m and the strength of the 1.6m H-bump are all larger than what expected by RGB alone, when diluted by the rest of the (featureless at these wavelengths) stellar phases (Fig.3e,f). This is the first time that multiple TiO absorptions concomitant with CN and C2 features are observed in a galaxy spectrum, pointing to the simultaneous presence of O-rich and C-rich cool stars. This is naturally expected for a broad distribution of stellar metallicities, which must exist in massive galaxies, such that the metal-poor stars evolve preferentially into C-rich TP-AGB stars, and the metal-rich ones into O-rich (M Type) TP-AGB stars[2]. Since V and Zr are only produced by s-process nucleosynthesis, those VO and ZrO features (Fig.2b), indicative of a recent third dredge-up episode, are expected only in the coolest TP-AGB stars[29, 31], demonstrating the prominence of the TP-AGB phase in the light emitted by this 1 quiescent galaxy.
@@
We subsequently performed spectral fitting on the spectrum of D36123, employing four models – BC03, C09, M13, and M05. These span the full range from TP-AGB poor to heavy (see Fig.2c,d and Table 1) and the first two are the standard for deriving the physical properties of high-redshift galaxies. We used delayed exponentially declining SFH with a large variety of timescales and stellar ages, interpolated over a grid of metallicities, and allowed for dust extinction. When convolving the models to match the resolution of observed spectra, we considered the instrumental resolution, the wavelength sampling of the spectral template models, and the estimated stellar velocity dispersion of the galaxy (see Methods). We find that our model fitting results are robust against (1) more complex SFHs allowing for bursts and sudden truncations; (2) adoption of non-parametric SFH obtained as linear combinations of models with different ages and metallicities; (3) alternative attenuation curves or without considering the reddening by dust; (4) variations to the aperture correction recipe; (5) modifications of the spectral extraction apertures over 3 or 5 pixels; (6) modelling of emission lines. The minimum reduced from each model is barely affected when these variations are considered.
@@
The “TP-AGB-mild/heavy” (i.e., M13 and M05) models can reproduce the largest fraction of the detected features, while the “TP-AGB-poor/light” predict much weaker spectral features in the NIR, as Fig.2c,d and Fig.3e,f illustrate. The TP-AGB-mild model, M13, provides the best overall fit and also identifies the majority of TP-AGB features (i.e., four TiO absorptions and most CN edges). However, there are strong spectral features that are clearly not present (or much weaker than in the observed spectrum) even in the M13 best-fit spectrum, like several TiO and C2 absorptions. The strong CN1.1m spectral edge is best reproduced by the M13/M05 models (Fig.3f). Interestingly, the MgII-band at 0.52m, part of Lick indexes[32] is much weaker in all models than in the data (Fig.3e), pointing to a strong contribution from TiO bands all the way to 0.5m, as expected for M-type stars (and, possibly, to [/Fe]-enhancement[33]). Overall, the M05 shows the worst fit due to clear discrepancy at 5000Å (see Extended Data Fig.4), but still has significantly more numerous matching TP-AGB features in the NIR than the C09 and BC03 models (Fig.2c,d), noticeably the long-sought C2 absorption around 1.75m. Moreover, when fitting the spectrum over the restricted 5000Å all models converge to a mass-weighted age of about 1 Gyr, which is indeed the age where TP-AGB features are expected to develop near peak[4, 11]. Therefore, we conclude that it is the treatment of the TP-AGB in the models that most affects the goodness of fit in our 1 quiescent galaxies. Even for the top performing model, M13, the minimum reduced 39 is large, indicating that also this model is lacking stellar ingredients that are present in the galaxy and are necessary.
@@
As mentioned, some observed features are also found in the coolest phase of (M-type) RGB stars (Fig.2b) and the red-supergiant (RSG) phase of massive stars, with ages around 10 Myr[13, 28, 31]. We could exclude that RSG stars contribute substantially due to the inferred mass-weighted age (1 Gyr) and negligible residual star formation rate (SFR) when fitting only the optical spectrum (see Table 1), implying a negligible contribution from short-lived massive stars. M-type RGB stars, on the other hand certainly contribute to the NIR spectrum at these ages, to a degree that is, however, model-dependent because the onset and development of the RGB phase depends on the overshooting assumed in stellar tracks[34, 4]. The full RGB contribution to the total light in the K-band around 1 Gyr is 20-30% according to M05 models and 10-20% for BC03-type models and the upper part of the RGB where M-giants form is a further fraction of this. Therefore, while contributing, M-type RGB stars are not expected to dominate the spectrum. This expectation is confirmed quantitatively by the measured indexes in Fig.3f, whose strengths are comparable or even larger than those measured on individual RGB stars, which are necessarily diluted by the composite populations in galaxies. Therefore, much strongly-featured TP-AGB stars are needed to match the observed strengths of the features in the galaxy.
@@
The best-fitting physical parameters presented in Table 1 exemplify the impact of the chosen model on the derived galaxy properties. For instance, comparing results between BC03 and M13 in the full rest-frame spectral range, there is only a 30% discrepancy in stellar mass, but the impact on ages is strong with stellar age difference 1 Gyr (the M13 age is almost three times lower than the BC03 age), despite the higher metallicity preferred by the BC03 fit. These age differences affect the derived formation epochs and evolutionary timescales of galaxies.
We also fitted additional models. Two entirely based on empirical stellar libraries (XSL[35] and E-MILES[36]; see Extended Data Fig.5a,b and Table 2 in Methods) without significant improvements over the quality of fit, or reproduction of features (Fig.3e,f). These models return old ages 1.5Gyr, with a sub-solar metallicity best-fit for E-MILES. XSL has closer EW indexes to what is observed in D36123 but performs less well on CN1.1m. The CB07 models (an updated version of BC03, see Methods) behave similarly to M05 in providing a worse global fit but better match to specific deep features (e.g., TiO@0.73m, Extended Data Fig.5c).